find center and radius of this circle
x^2+y^2+6x-4y+5=0

Answer by venugopalramana(3286)

find center and radius of this circle
x^2+y^2+6x-4y+5=0
(X^2+2*X*3+3^2)+(Y^2-2*Y*2+2^2)+5-3^2-2^2=0
(X+3)^2+(Y-2)^2=9+4-5=8
COMPARING WITH STD.EQN.OF IRCLE
(X-H)^2+(Y-K)^2=R^2...WHERE
CENTRE = (H,K)=(-3,2)
RADIUS = R = SQRT(8)=2SQRT(2)